OpenStudy (errinss):

What is the value of x? Round your answer to the nearest tenth. x^2 = 34

pooja195 (pooja195):

Square root both.

OpenStudy (errinss):

x=5.8309518948453004708741528775456

pooja195 (pooja195):

^leave the 34 as it is so that you don't need to work with that decimal

OpenStudy (errinss):

ok

pooja195 (pooja195):

x=?

OpenStudy (errinss):

\[\sqrt{x}=\sqrt{34}\]

pooja195 (pooja195):

since we square rooted we can get rid of the square root on the left (^2) cancels out.

pooja195 (pooja195):

Now, keep in mind there are 2 possible answers. We already have \[\sqrt{32 }\] what's the other answer?

OpenStudy (errinss):

wait isnt it 34?

pooja195 (pooja195):

^sorry i meant 34

OpenStudy (errinss):

lol its alright...

pooja195 (pooja195):

What would our answer be? :)

OpenStudy (cheez):

Ok here we go again. Your decimal was right - sqrt34 = now round it up like I taught you before: is 3 larger than 5?

OpenStudy (errinss):

oi

OpenStudy (cheez):

(I know that answer but I want you to figure it out) If 3 is larger than 5, then it becomes 5.9 which is your answer. If not, then 5.8.

pooja195 (pooja195):

@Cheez We don't use the decimal. It's easier to just leave it a square root.

pooja195 (pooja195):

5.9 is incorrect.

OpenStudy (errinss):

iye, oi, im confuffled....

OpenStudy (errinss):

*whelp*

OpenStudy (cheez):

Well I do and I'm 2 grades ahead of where I'm supposed to be in math. I'm a supposed to be a freshman but I'm in 10th grade with 11th grade algebra 2. I know what I'm doing.

OpenStudy (errinss):

*jaw drops*

OpenStudy (cheez):

Pooja I want errinss to figure it out for themselves. That way they learn something not just repeat words you hand out.

OpenStudy (errinss):

herself...

OpenStudy (cheez):

I thought so but you never know on the internet :P

OpenStudy (errinss):

X'D

OpenStudy (cheez):

5.8 is correct. You can even double check this. Instead of x, do 5.8 5.8^2 = ? Find ?

OpenStudy (errinss):

33.64...

OpenStudy (cheez):

Good. And since this isn't the full 5.8 number it is correct. you can round 33.64 up to 34.

jigglypuff314 (jigglypuff314):

Cheez has the right idea although I'm not sure where you got your "3" and "5" in your above explanation :)

OpenStudy (cheez):

need any more help?

pooja195 (pooja195):

@errinss We square rooted both sides which cancelled out the ^2 Since we are working with square roots it's best to leave the square root as it is. \[\huge~\rm~\bf~\sqrt{x^2}=\sqrt{34}\] You got to this part \[\huge~\rm~x=\sqrt{34}\] From the previous ones we have been doing, any square root will have two answers e.g: x^2=25 we would have 5 x 5=25 -5 x -5=25 we would have x=-5 or 5 Because if we ^2 a number by itself regardless if its negative or postive you will get the same solution.

OpenStudy (errinss):

lol ya

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Cheez Good. And since this isn't the full 5.8 number it is correct. you can round 33.64 up to 34. \(\color{#0cbb34}{\text{End of Quote}}\) You got the same answer.

pooja195 (pooja195):

But a different way.

OpenStudy (cheez):

And my way didn't include a bunch of mumbo jumbo. I'm teaching it the easier way.

OpenStudy (errinss):

im lost..... pooja are you trying to prove a point...im not seeing it

jigglypuff314 (jigglypuff314):

cheez, I never learned how to do it your way where did you get the "3" and the "5"?

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Cheez And my way didn't include a bunch of mumbo jumbo. I'm teaching it the easier way. \(\color{#0cbb34}{\text{End of Quote}}\) Alright, I'll let you take over then. :-) Sorry for confusing you @errinss

OpenStudy (cheez):

I meant * not 5 but I typed. Its 5.83 is 3 larger than 8.

pooja195 (pooja195):

3 is not larger than 8.

OpenStudy (errinss):

What is the value of y? 3y^2 − 6 = 42

jigglypuff314 (jigglypuff314):

I remember how to get that 25 < 34 < 36 but I'm not sure who you got that decimal after that.

jigglypuff314 (jigglypuff314):

I didn't see an explanation for that either

OpenStudy (errinss):

*lost in the clouds*

jigglypuff314 (jigglypuff314):

errinss, perhaps try asking your new question in a new question post :)

OpenStudy (cheez):

no need for a new post, but you can if you want. Add the six to both sides. Divide both sides by 3. then sqrt. What do you get?

jigglypuff314 (jigglypuff314):

nvm nvm I see it now I was thinking about doing it by hand (I'm old, that's how we did it back in the day) you had just been guiding on how to round, I see, sorry about the confusion .-.